# Deep Improvement Supervision

### v2

@@ 3 Different Faces of Latent Reasoning
Notation. Let the state be $\mathbf{s}=(\mathbf{x},\mathbf{z}_{L},\mathbf{z}_{H})$ at recursion step $t$, where $\mathbf{x}$ is the input, $\mathbf{z}_{L}$ is the latent reasoning scratchpad, and $\mathbf{z}_{H}$ is the current solution embedding. TRM applies a (parameter‑shared) update operator in two roles:
$\mathbf{z}_{L}^{n+1}=f_{\phi}^{(L)}(\mathbf{z}_{L}^{n},\mathbf{z}_{H}^{n},\mathbf{x}),\qquad\mathbf{z}_{H}^{n+1}=f_{\phi}^{(H)}(\mathbf{z}_{H}^{n},\mathbf{z}_{L}^{n+1}),$ (4)
where $f_{\phi}^{(L)}$ and $f_{\phi}^{(H)}$ are two instantiations that share parameters $\phi$ but differ in inputs. An output head $f_{O}$ maps $\mathbf{z}_{H}$ to a vector of logits. As established in §2.2, only $\mathbf{z}_{H}$ is decoded and trained (with cross‑entropy). We introduce a binary indicator $o\in\{0,1\}$ to distinguish two readout modes:
  • Reference ($o=0$): pre‑reasoning readout from $\mathbf{z}_{H}^{n}$, with $\ell_{u}:=f_{O}(\mathbf{z}_{H}^{n})$ is reference logits.
  • Conditional ($o=1$): post‑reasoning readout from $\mathbf{z}_{H}^{n+1}$ after attending to $\mathbf{z}_{L}^{n+1}$, with $\ell_{c}:=f_{O}(\mathbf{z}_{H}^{n+1})$ is conditional logits).
Throughout, we write $\pi_{u}=\mathrm{softmax}(\ell_{u})$, $\pi_{c}=\mathrm{softmax}(\ell_{c})$. Logits are understood as unnormalized log‑probabilities whose softmax defines the corresponding policy over actions/classes.
3.1 Latent Reasoning as Classifier‑Free Guidance
The Hierarchical Reasoning Model and the Tiny Recursive Model are based on iterative latent updates. As discussed in §2.1 and §2.2, the process is asymmetric: the reasoning state $\mathbf{z}_{L}$ is updated using $\mathbf{x}$ and $\mathbf{z}_{H}$, and this updated reasoning conditions the subsequent update of $\mathbf{z}_{H}$. At any recursion step $t$, this asymmetry yields two readouts:
$\ell_{u}=f_{O}(\mathbf{z}_{H}^{n})\text{(reference)},\qquad\ell_{c}=f_{O}(\mathbf{z}_{H}^{n+1})\text{(conditional)}.$
By interpolating in logit space we obtain guided logits and the corresponding guided policy:
$\ell_{w}\;=\;(1-w)\,\ell_{\mathrm{u}}\;+\;w\,\ell_{\mathrm{c}},\qquad\pi_{w}(\cdot\mid\mathbf{s})\;=\;\mathrm{softmax}(\ell_{w}).$ (5)
Equivalently, $\pi_{w}$ is a product of the two policies:
$\pi_{w}(a\mid\mathbf{s})\;\propto\;\pi_{u}(a\mid\mathbf{s})^{\,1-w}\;\pi_{c}(a\mid\mathbf{s})^{\,w}.$
Here, the residual $\Delta\ell:=\ell_{c}-\ell_{u}$ captures the direction of improvement suggested by latent reasoning, while $w$ acts as a guidance scale (or reasoning depth).
Practical note. In standard TRM inference, decoding uses the conditional path $\mathbf{z}_{H}^{n+1}$ (i.e., $w=1$). Eq. (5) exposes $w$ as a controllable knob at test time, allowing amplification or attenuation of the reasoning signal. [^1] Classifier‑Free Guidance in diffusion mixes scores (gradients of log density). Here we mix logits. The algebraic product‑of‑experts identity above holds because logits exponentiate to probabilities under softmax.
3.2 Latent Reasoning as Implicit Policy Improvement
Section 3.1 shows that TRM’s two readouts can be mixed as in CFG. We now show that the same mechanism implements a controllable policy improvement operator. Consider sampling from a product policy:
$\pi_{\text{new}}(y\mid\mathbf{s})\;\propto\;\underbrace{\hat{\pi}(y\mid\mathbf{s})}_{\text{Reference policy}}\cdot\underbrace{p(o=1\mid\mathbf{s},y)}_{\text{Optimality likelihood}},$ (6)
where $p(o=1\mid\mathbf{s},y)$ measures how likely action $y$ leads to an optimal outcome (or, in our approximation below, to an improved conditional distribution).
Guidance‑based RL (frans2025diffusion) shows that such product policies improve performance under mild conditions: if $p(o=1\mid\mathbf{s},y)$ increases monotonically with the advantage $A_{\hat{\pi}}(\mathbf{s},y)$, then the product policy has higher expected return than $\hat{\pi}$. Introducing a guidance scale $w\geq 0$ and applying Bayes’ rule,
$\log p(o\mid\mathbf{s},y)=\log\hat{\pi}(y\mid\mathbf{s},o)-\log\hat{\pi}(y\mid\mathbf{s})+\log p(o\mid\mathbf{s}),$
where $\log p(o\mid\mathbf{s})$ is independent of $y$, we obtain
$\log\pi_{\text{new}}(y\mid\mathbf{s})$ $\propto\log\hat{\pi}(y\mid\mathbf{s})+$
$\;w\big(\log\hat{\pi}(y\mid\mathbf{s},o)-\log\hat{\pi}(y\mid\mathbf{s})\big)$
$\;\equiv\;(1-w)\,\underbrace{\log\hat{\pi}(y\mid\mathbf{s})}_{\text{``old''}}+w\,\underbrace{\log\hat{\pi}(y\mid\mathbf{s},o)}_{\text{``new''}}.$
Exponentiating and normalizing yields the exact identity
$\boxed{\;\pi_{w}(y\mid\mathbf{s})\;\propto\;\hat{\pi}(y\mid\mathbf{s})\cdot\left(\frac{\hat{\pi}(y\mid\mathbf{s},o=1)}{\hat{\pi}(y\mid\mathbf{s})}\right)^{\!w}\;}$ (7)
which matches the logit interpolation in Eq. (5) when $\ell_{u}$ and $\ell_{c}$ are the (unnormalized) log‑probabilities of $\hat{\pi}(\cdot\mid\mathbf{s})$ and $\hat{\pi}(\cdot\mid\mathbf{s},o=1)$, respectively.
By training the conditional path ($\mathbf{z}_{H}^{n+1}$, hence $\ell_{c}$) toward the correct answer while the reference path ($\mathbf{z}_{H}^{n}$, hence $\ell_{u}$) encodes the pre‑reasoning baseline, the residual $\Delta\ell=\ell_{c}-\ell_{u}$ functions as a log‑likelihood ratio—an advantage‑like update that reweights actions the reasoning step deems better. Thus TRM performs a step of policy iteration at each forward pass without explicitly estimating a value function.
@@ 4 Deep Improvement Supervision
While §3.2 establishes that TRMs implicitly perform policy improvement, relying solely on terminal supervision creates a difficult optimization landscape: the model must discover a sequence of latent updates that yields a correct final answer, inferring the direction of improvement from a sparse, high‑variance terminal signal. To address this, we introduce Deep Improvement Supervision (DIS). We first formalize the condition under which the asymmetric readouts improve predictions—Advantage Margin—and then show how DIS directly optimizes this condition by turning “reasoning” into a concrete, step‑wise supervised objective.
4.1 The Mechanics of Improvement
At recursion step $t$, the TRM produces two sets of logits:
  • Reference logits $\ell_{u}$: from $\mathbf{z}_{H}^{n}$, representing the policy before the reasoning step.
  • Conditional logits $\ell_{c}$: from $\mathbf{z}_{H}^{n+1}$, after attending to $\mathbf{z}_{L}^{n+1}$.
Define the residual $\Delta\ell(a)=\ell_{c}(a)-\ell_{u}(a)$. With guidance weight $w$, $\pi_{w}=\mathrm{softmax}(\ell_{u}+w\Delta\ell)$. Let $y^{\star}$ denote the correct class at the current decoding position. We ask: under what condition does increasing $w$ improve $\pi_{w}(y^{\star})$?
Proposition 4.1 (Advantage Margin Condition).
Let $\mathcal{L}(w)=-\log\pi_{w}(y^{\star})$ be the cross‑entropy loss for the correct class. Then $\frac{d}{dw}\mathcal{L}(w)<0$ (i.e., the loss strictly decreases as $w$ increases) if and only if
$\underbrace{\Delta\ell[y^{\star}]}_{\text{Boost to correct class}}\;>\;\underbrace{\mathbb{E}_{a\sim\pi_{w}}\big[\Delta\ell[a]\big]}_{\text{Policy‑weighted average boost}}.$ (8)
Moreover, $\frac{d}{dw}\mathcal{L}(w)=\mathbb{E}_{a\sim\pi_{w}}[\Delta\ell[a]]-\Delta\ell[y^{\star}]$ and $\frac{d^{2}}{dw^{2}}\mathcal{L}(w)=\mathrm{Var}_{a\sim\pi_{w}}[\Delta\ell[a]]\geq 0$.
See Appendix A for details. Intuitively, reasoning helps when the residual gives the correct prediction a larger boost than it gives a typical alternative, in other words, when Advantage Margin is positive.
4.2 From Implicit Discovery to Explicit Supervision
In standard TRM training, the model attempts to satisfy Proposition 4.1 implicitly. DIS explicitly enforces a positive Advantage Margin at every recursive step by constructing intermediate targets that contract toward the solution.
Let $\mathbf{x}$ be the input and $\mathbf{y}^{\star}$ the ground‑truth output sequence. Let $\Phi$ be a target generator (e.g., a discrete diffusion‑style schedule) that produces a sequence $\{\mathbf{y}^{\dagger}_{s}\}_{s=0}^{N_{\text{sup}}}$ with $\mathbf{y}^{\dagger}_{N_{\text{sup}}}=\mathbf{y}^{\star}$ and each $\mathbf{y}^{\dagger}_{s}$ strictly closer to $\mathbf{y}^{\star}$ than $\mathbf{y}^{\dagger}_{s-1}$ under a fixed metric, for example Hamming distance.
Ideally, we want the reference logits at step $s$ to match the previous target $\mathbf{y}^{\dagger}_{s-1}$ and the conditional logits to match the improved target $\mathbf{y}^{\dagger}_{s}$, leading to a double‑loss objective:
$\mathcal{L}_{\text{Dual}}=\sum_{s=1}^{N_{\text{sup}}}\Big[\underbrace{\mathrm{CE}(\ell_{u}^{s},\mathbf{y}^{\dagger}_{s-1})}_{\text{Anchor previous}}\;+\;\underbrace{\mathrm{CE}(\ell_{c}^{s},\mathbf{y}^{\dagger}_{s})}_{\text{Predict current}}\Big].$ (9)
However, in a recursive architecture, this is redundant: the input to step $s$ (which produces $\ell_{u}^{s}$) is a deterministic function of the output of step s1s-1 (which produced $\ell_{c}^{s-1}$). Since step s1s-1 is trained toward $\mathbf{y}^{\dagger}_{s-1}$, $\ell_{u}^{s}$ is implicitly anchored. We therefore use the single‑loss objective
@@ 4.2 From Implicit Discovery to Explicit Supervision
Because optimization is sequential, minimizing Eq. (10) is (under teacher‑forced recursion) inductively equivalent to minimizing Eq. (9): forcing $\ell_{c}^{s-1}\!\to\!\mathbf{y}^{\dagger}_{s-1}$ ensures $\ell_{u}^{s}$ starts near $\mathbf{y}^{\dagger}_{s-1}$.
Training $\ell_{c}$ on $\mathbf{y}^{\dagger}_{s}$ while $\ell_{u}$ is anchored to $\mathbf{y}^{\dagger}_{s-1}$ forces the residual $\Delta\ell=\ell_{c}-\ell_{u}$ to encode the transition from $\mathbf{y}^{\dagger}_{s-1}$ to $\mathbf{y}^{\dagger}_{s}$. This yields our second formal result.
Proposition 4.2 (Guaranteed Improvement).
Assume the target generator $\Phi$ provides strictly improving targets with respect to a fixed scoring distribution $P$, i.e., $\log\tfrac{P(\mathbf{y}^{\dagger}_{s})}{P(\mathbf{y}^{\dagger}_{s-1})}>0$ for all $s$. Then minimizing the sequential loss $\mathcal{L}_{\text{DIS}}$ drives the expected Advantage Margin to be positive:
$\mathbb{E}\Big[\Delta\ell[y^{\star}]-\mathbb{E}_{a\sim\pi_{w}}\Delta\ell[a]\Big]\;>\;0,$
where the outer expectation is over data (and model stochasticity) and the inner expectation is with respect to $\pi_{w}$ for any fixed $w\geq 0$.
See Appendix A for details. Standard training hopes the model learns to improve; DIS forces each residual vector $\Delta\ell$ to point toward the solution. Consequently, DIS turns TRM from a black‑box recurrent model into a verifiable iterative refinement procedure in which every forward pass is supervised to act as a precise policy‑improvement step.

### v3

@@ 3 Formalization of Latent Reasoning as Policy Improvement
where $x$ is the embedded input, $z_{L}^{t}$ is a latent “scratchpad”, and $z_{H}^{t}$ is the current solution embedding. A parameter-shared backbone is used in two roles:
$z_{L}^{t+1}=f_{\phi}^{(L)}(x,z_{H}^{t},z_{L}^{t}),\qquad z_{H}^{t+1}=f_{\phi}^{(H)}(z_{H}^{t},z_{L}^{t+1}),$ (6)
and an output head $f_{O}$ maps $z_{H}$ to logits over the next token/action set $\mathcal{A}$. Only $z_{H}$ is directly decoded and trained with cross-entropy.
3.1 Two policies per recursion answer
Each recursion step produces two output distributions at essentially no extra cost: one before the latent update and one after it.
Reference policy (pre-reasoning). Decode before updating:
$\ell_{u}^{t}\;:=\;f_{O}(z_{H}^{t}),\qquad\hat{\pi}_{t}(a\mid s_{t})\;:=\;\mathrm{softmax}(\ell_{u}^{t})[a].$ (7)
Improved policy (post-reasoning). Decode after one latent-reasoning update:
$\ell_{c}^{t}\;:=\;f_{O}(z_{H}^{t+1}),\qquad\pi_{t}^{+}(a\mid s_{t})\;:=\;\mathrm{softmax}(\ell_{c}^{t})[a].$ (8)
Intuitively, $\hat{\pi}_{t}$ is the model’s first guess at step $t$, while $\pi_{t}^{+}$ is the refined distribution after one reasoning step.
We now introduce the same object used in policy-improvement analyses: an optimality (or improvement) variable. Let $o\in\{0,1\}$ indicate whether an action $a$ is “good” for state $s$. (For supervised next-token prediction, the simplest choice is $o=1$ iff $a$ equals the ground-truth token; more general choices can encode task-level reward or self-consistency.) A standard KL-regularized improvement step produces a new policy by reweighting a reference policy $\hat{\pi}$ using an optimality likelihood $p(o=1\mid s,a)$:
$\pi_{w}(a\mid s)\;\propto\;\hat{\pi}(a\mid s)\,p(o=1\mid s,a)^{\,w},$ (9)
where $w\geq 0$ controls the step size: larger $w$ places more weight on actions deemed optimal. Equation (9) is exactly the multiplicative form of an exponentiated-advantage update; explicitly, if we define the improvement score
$A(s,a)\;:=\;\log p(o=1\mid s,a),$ (10)
then (9) is equivalent to
$\pi_{w}(a\mid s)\;=\;\frac{\hat{\pi}(a\mid s)\,\exp\!\big(w\,A(s,a)\big)}{\mathbb{E}_{a^{\prime}\sim\hat{\pi}(\cdot\mid s)}\!\left[\exp\!\big(w\,A(s,a^{\prime})\big)\right]}.$ (11)
This view matches a general policy-improvement theorem for product policies. If we form a new policy by multiplying a reference policy by a non-negative, monotonically increasing function of advantage, the result is guaranteed to improve expected return.Let $f:\mathbb{R}\to\mathbb{R}_{\geq 0}$ be non-negative and monotonically increasing. If $A^{\hat{\pi}}(s,a)$ denotes the advantage under $\hat{\pi}$, then the product policy
$\pi(a\mid s)\;\propto\;\hat{\pi}(a\mid s)\,f\!\big(A^{\hat{\pi}}(s,a)\big)$ (12)
is an improvement over $\hat{\pi}$. Please, see Remark 1 and Theorem 1 (Frans et al., 2025). We can say that, (11) is a special case of (12) with $f(u)=\exp(wu)$.
The only missing ingredient is: what is $p(o=1\mid s,a)$ for latent reasoning? In our setting, we will recover an implicit improvement factor $p(o=1\mid s,a)$ from the TRM’s own post-reasoning policy, yielding the same product-policy form, but without explicitly learning a value function.
3.2 Improvement from a conditional policy
Here is the key observation: TRM already gives us a natural candidate for the optimality-conditioned policy. Interpret the post-reasoning distribution as
$\pi_{t}^{+}(a\mid s_{t})\;\approx\;p(a\mid s_{t},o=1),$ (13)
i.e., “the policy after one reasoning step behaves like the policy conditioned on the event that the answer is improved.”
Under this interpretation we can recover an implicit optimality model by Bayes’ rule:
$p(o=1\mid s,a)\;=\;\frac{p(a\mid s,o=1)\,p(o=1\mid s)}{p(a\mid s)}.$
Substituting $p(a\mid s,o=1)\approx\pi_{t}^{+}(a\mid s_{t})$ and $p(a\mid s)\approx\hat{\pi}_{t}(a\mid s_{t})$ yields
$p(o=1\mid s_{t},a)\;\propto\;\frac{\pi_{t}^{+}(a\mid s_{t})}{\hat{\pi}_{t}(a\mid s_{t})},$ (14)
where the proportionality constant $p(o=1\mid s_{t})$ does not depend on $a$. Plugging (14) into the generic improvement step (9) gives
$\pi_{t,w}(a\mid s_{t})$ $\propto\hat{\pi}_{t}(a\mid s_{t})\left(\frac{\pi_{t}^{+}(a\mid s_{t})}{\hat{\pi}_{t}(a\mid s_{t})}\right)^{w}$ (15)
$\propto\hat{\pi}_{t}(a\mid s_{t})^{\,1-w}\,\pi_{t}^{+}(a\mid s_{t})^{\,w}.$ (16)
Equation (16) is the policy-improvement analogue of the “Bayes inversion” step: it rewrites a multiplicative optimality reweighting in terms of a pair of policies (unconditional vs. optimality-conditioned). Concretely, latent reasoning provides both factors $\hat{\pi}_{t}$ and $\pi_{t}^{+}$, and therefore implicitly defines the entire one-step improvement family $\{\pi_{t,w}\}_{w\geq 0}$.
From (14), the improvement score is the log-ratio
$A_{t}(s_{t},a)$ $:=\log\pi_{t}^{+}(a\mid s_{t})-\log\hat{\pi}_{t}(a\mid s_{t})$ (17)
$\equiv\log p(o=1\mid s_{t},a)\;+\;\mathrm{const}(s_{t}),$
i.e., up to a state-dependent constant, the TRM step computes an advantage-like signal: it measures how much the reasoning step increases the relative probability of action $a$.
3.3 A supervised improvement criterion
For next-token prediction, let $y^{\star}$ be the ground-truth token at the current position. Define the cross-entropy under the improved policy family (16):
$L_{t}(w)\;:=\;-\log\pi_{t,w}(y^{\star}\mid s_{t}).$
Writing $\pi_{t,w}(\cdot\mid s_{t})=\mathrm{softmax}\!\big((1-w)\log\hat{\pi}_{t}+w\log\pi_{t}^{+}\big)$, a direct derivative gives
$\frac{d}{dw}L_{t}(w)\;=\;\mathbb{E}_{a\sim\pi_{t,w}(\cdot\mid s_{t})}\!\big[A_{t}(s_{t},a)\big]\;-\;A_{t}(s_{t},y^{\star}),$
and
$\frac{d^{2}}{dw^{2}}L_{t}(w)\;=\;\mathrm{Var}_{a\sim\pi_{t,w}(\cdot\mid s_{t})}\!\big[A_{t}(s_{t},a)\big]\;\geq\;0.$
Thus increasing $w$ (trusting the reasoning step more strongly) decreases the loss if and only if the ground-truth action has above-average improvement score:
$A_{t}(s_{t},y^{\star})\;>\;\mathbb{E}_{a\sim\pi_{t,w}(\cdot\mid s_{t})}\!\big[A_{t}(s_{t},a)\big].$ (18)
Equation (18) is a precise, testable condition for when a latent-reasoning update is helpful: the step improves prediction exactly when it preferentially increases the relative log-probability of the correct token compared to typical alternatives under the current policy.
3.4 Latent recursion as repeated policy improvement
Putting the pieces together, one TRM recursion step produces: (i) a reference policy $\hat{\pi}_{t}$, (ii) an approximate optimality-conditioned policy $\pi_{t}^{+}$, and therefore (iii) an implicit improvement model $p(o=1\mid s_{t},a)$ via Bayes inversion (14). The family $\pi_{t,w}$ in (16) is exactly the corresponding policy-improvement update, with an advantage-like signal given by the log-ratio.
The recursion $t\mapsto t+1$ can therefore be read as iterating a learned policy improvement operator on a fixed input: each latent step proposes a new distribution over actions and simultaneously provides the internal signal that justifies reweighting the previous distribution. The remaining question is not whether recursion converges, but whether training shapes $A_{t}(s,a)$ to be consistently aligned with task success—so that each step performs non-dead, directed improvement rather than redundant computation.
@@ 4 Deep Improvement Supervision
Section 3 reframes a latent-reasoning step as a policy-improvement operator: each recursion produces a reference policy $\hat{\pi}_{t}(\cdot\mid s_{t})$ and a post-reasoning policy $\pi_{t}^{+}(\cdot\mid s_{t})$, and their log-ratio $A_{t}(s_{t},a)=\log\pi_{t}^{+}(a\mid s_{t})-\log\hat{\pi}_{t}(a\mid s_{t})$ acts as an advantage-like improvement signal. The analysis yielded a concrete Advantage Margin condition: a latent update helps exactly when the ground-truth action has above-average improvement score under the interpolated policy family $\pi_{t,w}$ (Eq. (18)).
[…]
Deep Improvement Supervision (DIS) operationalizes this idea by providing explicit stepwise improvement targets and supervising the post-reasoning readout at every recursion step so that each latent update realizes a directed improvement. Let $\mathbf{x}$ be the input and $\mathbf{y}^{\star}$ the ground-truth output sequence. Let $\Phi$ be a target generator that produces a sequence $\{\mathbf{y}^{\dagger}_{s}\}_{s=0}^{N_{\text{sup}}}$ with $\mathbf{y}^{\dagger}_{N_{\text{sup}}}=\mathbf{y}^{\star}$ and $\mathbf{y}^{\dagger}_{s}$ strictly closer to $\mathbf{y}^{\star}$ than $\mathbf{y}^{\dagger}_{s-1}$ under a fixed discrepancy metric (e.g., expected Hamming distance).
We interpret the supervision step $s$ as enforcing a single policy-improvement move: the pre-reasoning policy $\hat{\pi}_{s}$ should behave like a policy centered on $\mathbf{y}^{\dagger}_{s-1}$, while the post-reasoning policy $\pi_{s}^{+}$ should behave like the improved (optimality-conditioned) policy centered on $\mathbf{y}^{\dagger}_{s}$. The most literal objective is therefore a dual cross-entropy:
$\mathcal{L}_{\text{Dual}}=\sum_{s=1}^{N_{\text{sup}}}\Big[\underbrace{\mathrm{CE}(\ell_{u}^{s},\mathbf{y}^{\dagger}_{s-1})}_{\text{Anchor reference}}\;+\;\underbrace{\mathrm{CE}(\ell_{c}^{s},\mathbf{y}^{\dagger}_{s})}_{\text{Supervise improvement}}\Big].$ (19)
In a recursive architecture, the input to step $s$ (which produces $\ell_{u}^{s}$) is a deterministic function of the post-reasoning state of the previous step (which produces $\ell_{c}^{s-1}$). Under teacher-forced recursion, training $\ell_{c}^{s-1}$ toward $\mathbf{y}^{\dagger}_{s-1}$ implicitly anchors $\ell_{u}^{s}$ near the previous target. Thus we use the single-loss
@@
Inductively, minimizing Eq. (20) enforces the same stepwise anchoring as Eq. (19): forcing $\ell_{c}^{s-1}\!\to\!\mathbf{y}^{\dagger}_{s-1}$ ensures $\ell_{u}^{s}$ starts near $\mathbf{y}^{\dagger}_{s-1}$.
4.1 DIS as direct optimization of the Advantage Margin
Training $\ell_{c}$ on $\mathbf{y}^{\dagger}_{s}$ while $\ell_{u}$ is anchored to $\mathbf{y}^{\dagger}_{s-1}$ forces the residual $\Delta\ell^{s}=\ell_{c}^{s}-\ell_{u}^{s}$ to encode the incremental improvement direction from $\mathbf{y}^{\dagger}_{s-1}$ to $\mathbf{y}^{\dagger}_{s}$. Equivalently, in probability space, DIS shapes the log-ratio advantage $A_{s}(a)=\log\pi_{s}^{+}(a\mid s)-\log\hat{\pi}_{s}(a\mid s)$ assign higher-than-average advantage to the next target token(s), aligning with the theoretical Advantage Margin condition (Eq. (18)). This converts the previous section’s existence/diagnostic statement (“improvement occurs iff the margin is positive”) into a constructive training rule: make the margin positive at every step by supervising the post-reasoning policy on a strictly improving target sequence.
Proposition 4.1 (Stepwise Advantage-Margin Alignment).
Assume that the target generator $\Phi$ provides strictly improving targets with respect to a fixed scoring distribution $P$, i.e., $\log\tfrac{P(\mathbf{y}^{\dagger}_{s})}{P(\mathbf{y}^{\dagger}_{s-1})}>0$ for all $s$. Then minimizing the sequential loss $\mathcal{L}_{\text{DIS}}$ drives the expected Advantage Margin to be positive:
$\mathbb{E}\Big[\Delta\ell[\mathbf{y}^{\star}]-\mathbb{E}_{a\sim\pi_{w}}\Delta\ell[a]\Big]\;>\;0,$
where the outer expectation is over the data (and model stochasticity) and the inner expectation is with respect to $\pi_{w}$ for any fixed $w\geq 0$.
See Appendix B for more details. In short: standard training relies on the hope that recursion happens to learn a monotone $f(A)$-style reweighting. DIS enforces it by construction, by making each post-reasoning policy $\pi_{s}^{+}$ predict an explicitly improved target relative to the implicitly anchored reference policy $\hat{\pi}_{s}$. The result is a verifiable iterative refinement procedure in which each forward step is trained to behave like a concrete policy-improvement move.